OpenStudy (anonymous):

HELP PLEASE!!! Consider the two vectors A = 3 i − 3 j and B = −i − 6 j . (a) Calculate A + B (b) Calculate A − B (c) Calculate | A + B| (d) Calculate | A - B| (e) Calculate the directions of A + B and A − B. A + B = ° (counterclockwise from the +x axis) A− B = ° (counterclockwise from the +x axis)

OpenStudy (anonymous):

A = 3 i − 3 j and B = −i − 6 j A+B = ( 3i -3j ) +( -i-6j ) =3i-i -3j-6j do the last step

OpenStudy (anonymous):

i got -6 for B and that was wrong

OpenStudy (anonymous):

A+B = 2i -9j is the answer

OpenStudy (anonymous):

now A-B = ( 3i -3j ) - ( -i-6j ) = 3i +i -3j+6j

OpenStudy (anonymous):

for A-B i got 6 - 3 the 3 was right and 6 was wrong

OpenStudy (anonymous):

yes its not 6 its -9

OpenStudy (anonymous):

that was wrong

OpenStudy (anonymous):

so i told you the answer take that

OpenStudy (anonymous):

thats not for A - B. so stp being rude

OpenStudy (amistre64):

i prefer a vertical setup A: 3 i - 3 j B: - i - 6 j adding it is simple enough, just work the columns as is A: 3 i - 3 j B: - i - 6 j ----------- 2i - 9j ------------------------------------------------- subtraction is just a matter of altering it as needed and working the columns; A = 3 i - 3 j -B = i + 6 j ------------- 4i + 3j

OpenStudy (amistre64):

c and d are asking for the length of the vectors from the first questions